Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-8x – 4y = -68
x + 3y = 16

Answer by mananth(16946)

-8x-4y=-68 .............1
3y=16-x ---------2
/3
y=5.33-0.33x

Substitute y in (1)
-8x-4(5.33-0.33 x)=-68
-8x-21.33+1.33x=-68
-8x+1.33x=-68+ 21.33
-6.67 x = -46.67
/-6.67
x=7
Plug the value of x in (1)
-8x-4y =-68
-8*7-4y=-68
-56-4y = -68
-4y=-12
/-4
y=3